What is -3/10+a=2/3

Found 2 solutions by tommyt3rd, Alan3354:
Answer by tommyt3rd(5050)

-3/10+a=2/3
a=2/3+3/10
a=20/30+9/30
a=29/30

Answer by Alan3354(69443)

What is -3/10+a=2/3
-----------
It's an equation.